# Cognitive Foundations for Reasoning and Their Manifestation in LLMs

### v1

@@ Abstract
Large language models solve complex problems yet fail on simpler variants, suggesting they achieve correct outputs through mechanisms fundamentally different from human reasoning. We synthesize cognitive science research into a taxonomy of 28 cognitive elements spanning computational constraints, meta-cognitive controls, knowledge representations, and transformation operations, then analyze their behavioral manifestations in reasoning traces. We propose a fine-grained cognitive evaluation framework and conduct the first large-scale analysis of 170K traces from 17 models across text, vision, and audio modalities, alongside 54 human think-aloud traces, which we make publicly available. Our analysis reveals systematic structural differences: humans employ hierarchical nesting and meta-cognitive monitoring while models rely on shallow forward chaining, with divergence most pronounced on ill-structured problems. Meta-analysis of 1,598 LLM reasoning papers reveals the research community concentrates on easily quantifiable behaviors (sequential organization: 55%, decomposition: 60%) while neglecting meta-cognitive controls (self-awareness: 16%, evaluation: 8%) that correlate with success. Models possess behavioral repertoires associated with success but fail to deploy them spontaneously. Leveraging these patterns, we develop test-time reasoning guidance that automatically scaffold successful structures, improving performance by up to 60% on complex problems. By bridging cognitive science and LLM research, we establish a foundation for developing models that reason through principled cognitive mechanisms rather than brittle spurious reasoning shortcuts or memorization, opening new directions for both improving model capabilities and testing theories of human cognition at scale.
@@ 1 Introduction
Humans are capable of extrapolating from their existing knowledge to unfamiliar scenarios and generating new knowledge—a process that constitutes reasoning (lombrozo2024learning). In contrast, large language models (LLMs) exhibit a form of cognitive dissonance unprecedented in human cognition: they master complex skills (chervonyi2025gold; jiang2024survey) while lacking simpler prerequisite ones (mccoy2024embers; dziri2023faith; berglund2023reversal), and solve challenging problems while failing on trivial variants (shao2025spurious; li2025personalized). This dissociation between high benchmark performance and the lack of generalization suggests that models may be achieving correct outputs through mechanisms fundamentally different from the robust cognitive structures underlying human reasoning. Current training and evaluation paradigms reward reasoning outcomes without examining the cognitive processes that produce them (lambert2024tulu3), and therefore cannot distinguish between genuine reasoning and memorization (wu2025reasoning). This creates a measurement crisis: we lack the conceptual framework to characterize what reasoning behaviors should manifest in models and the empirical methods to assess whether they do.
@@
To understand how cognitive elements manifest in reasoning behaviors, consider a child playing with LEGO blocks. Figure  1 illustrates how the child orchestrates multiple cognitive elements: establishing goals, decomposing the task into parts, verifying connections, monitoring progress, and reformulating failed approaches. This flexible coordination of diverse cognitive processes characterizes reasoning. However, contemporary research on LLM reasoning only attempts scattered investigations of specific behaviors such as verification (gandhi2025cognitive), decomposition (xu2025large), and self-monitoring (marjanovic2025deepseek). Our meta-analysis of 1,598 arXiv LLM reasoning papers on LLM reasoning reveals this fragmentation quantitatively (Figure 2): 55% examine sequential organization (e.g. reasoning step-by-step), while only 8% study backward chaining (e.g. reverse engineering the solution from the answer). This concentration on easily quantifiable behaviors creates a striking asymmetry when compared with human reasoning research, which encompasses diverse phenomena, including analogical transfer (gentner1983structure; holyoak1989analogical), causal inference (sloman2009causal; gopnik2004theory), representational flexibility (ohlsson1992information; knoblich1999constraint), and meta-cognitive control (nelson1990metamemory; flavell1979metacognition). Each of these captures different facets of how humans navigate uncertainty, integrate knowledge, and adapt to novel situations. Without a comprehensive framework spanning this range of cognitive elements, we risk optimizing what we measure rather than what matters, potentially mistaking narrow competence for broad reasoning capability.
@@
Figure 1: An example of reasoning behaviors present in a reasoning trace for building a LEGO spaceship. We organize the different behaviors along the four dimensions of our taxonomy, as shown in Table 2.
@@
Figure 2: Distribution of cognitive element presence across 1,598 arXiv LLM Reasoning papers. Partially present indicates that there is evidence that the behavior was considered in the design (motivation, method, evaluation) of the paper but was not the primary focus. Present indicates that there is evidence that the behavior was a conscious and significant design decision. Details provided in Section 5.
@@
To address this gap, we introduce a unified taxonomy of cognitive foundations by synthesizing established theories of problem-solving, mental representation, and meta-cognition (sweller1988cognitive; johnson1983mental; fodor1975language). The taxonomy includes four dimensions: 1) Reasoning invariants capture fundamental properties that must hold for valid reasoning, including logical coherence (not simultaneously believing "the design is stable" and "the design will collapse"), compositionality (understanding "red LEGO cockpit" by combining concepts of color, material, and function). 2) Meta-cognitive controls encompass the executive functions that select and monitor reasoning strategies, such as recognizing when you lack necessary pieces (self-awareness), deciding whether to plan the entire design upfront versus building exploratively (strategy selection). 3) Reasoning representations desrcibe how knwoledge is organized: hierarchically (decomposing "spaceship" into "body," "wings," "cockpit"), causally (understanding that inadequate support causes structural collapse), or spatially (tracking how pieces connect in 3D space). 4) Reasoning operations specify the procedures that construct and transform these representations, such as verifying each connection works, backtracking when a wing design fails, or abstracting the principle that heavier components need more support points. Organizing these theories through Marr’s (marr1982vision) levels of analysis (focusing on computational and algorithmic levels) yields 28 specific cognitive elements spanning the space of human reasoning capabilities (Table 2). This taxonomy represents a theoretical contribution: we synthesize existing cognitive science literature into an analytical vocabulary for studying machine reasoning, providing a bridge between human cognitive research and LLM evaluation.
@@
With this taxonomy, we conduct the first large-scale empirical comparison of cognitive elements in human versus LLM reasoning across diverse problem types (jonassen2000toward). We analyze 170K reasoning traces from 17 models across text, vision, and audio modalities, complemented by 54 human think-aloud reasoning traces (Table 2). Using fine-grained span-level annotation validated by human evaluation, we identify which of the 28 cognitive elements appear in each trace, where they occur, and how they are sequenced in thought. Our analysis reveals fundamental differences in reasoning presence and structures. We apply our analysis framework to distinguish between elements that are frequently exhibited by models and those that are most strongly correlated with correct outcomes. Through our analysis, we reveal that models consistently employ behaviors which are not the most conducive to success. We specifically observe that as problems become more ill-defined and non-verifiable (e.g., open-ended dilemmas, diagnosing and solving multifaceted problems), models narrow their selection of elements to rigid strategies despite a broader, more diverse usage of cognitive elements being empirically shown as a more effective strategy. Furthermore, we devise a novel structural representation to encode the structure of cognitive elements within reasoning traces and find that models frequently choose reasoning structures which differ from successful structure. Finally, we apply this same analysis framework to comparing humans and LLMs. We observe that humans employ more frequent abstract and conceptual-level processing, while LLMs rely more heavily on shallow sequential forward chaining with limited corrective structures based on problem type and structure.
@@
Leveraging these behavioral correlations, we introduce test-time reasoning guidance as a targeted intervention to improve model performance by explicitly scaffolding cognitive patterns predictive of reasoning success. For instance, on dilemma problems where successful DeepSeek-R1 traces exhibit the behavioral sequence of self-awareness followed by hierarchical representation construction and then problem decomposition, we prompt the model to follow this structure. Evaluation demonstrates that test-time reasoning guidance improves performance by up to 66.7% on ill-structured problems while maintaining baseline performance on well-structured ones. This improvement confirms that models possess reasoning capabilities not spontaneously expressed, and that understanding cognitive behavioral patterns can inform more effective model interaction strategies.
@@
This work establishes a taxonomy of cognitive foundations that bridges human reasoning research and LLM evaluation, providing a unified vocabulary for characterizing reasoning processes beyond performance metrics. By synthesizing cognitive science theories through Marr’s levels of analysis, we identify 28 reasoning behaviors spanning computational constraints, meta-cognitive controls, knowledge representations, and transformation operations. We conduct the first large-scale empirical analysis across 170K model and human reasoning traces [^1] Publicly released at https://huggingface.co/collections/stellalisy/cognitive-foundations, revealing systematic structural differences: humans employ hierarchical nesting and meta-cognitive monitoring while models rely on shallow sequential forward chaining, with divergence most pronounced on ill-structured problems. Leveraging these patterns, we introduce cognitively-grounded reasoning guidance to the model that automatically scaffolds successful reasoning structures, improving performance by up to 60% on complex problems and confirming that models possess latent capabilities not spontaneously expressed. Our analysis of 1,598 LLM reasoning papers reveals critical between research emphasis and behavioral requirements. This framework enables researchers to ask targeted questions about LLM reasoning: which behaviors emerge from scale and which require architectural innovation, do different training paradigms produce different cognitive profiles, and how do reasoning structures transfer across domains and modalities? These questions become answerable when we have a shared vocabulary for characterizing reasoning processes. Our taxonomy and annotation methodology provide that vocabulary, grounded in decades of cognitive science research and validated through large-scale empirical analysis of both human and machine reasoning traces.
@@ 2 Formalizing Cognitive Foundations for Reasoning
Table 1: Taxonomy of cognitive reasoning behaviors, organized along two main dimensions: Cognitive Properties (stable internal constraints) and Cognitive Capabilities (deployable context-sensitive abilities).
@@
Modification heightAdaptive detail mgmt. Adjust granularity based on task demands (Rosch1978-ROSPOC-8).
@@
Understanding the extent to which language models reason requires a theory of what reasoning is. Without principled criteria distinguishing reasoning from pattern matching, evaluation isn’t possible. Decades of research in Cognitive Science on problem-solving, mental representation, and metacognition provide a powerful starting point. The cognitive revolution of the 1950s introduced a powerful framework: the mind as an information-processing system. Fodor’s (fodor1975language) Language of Thought Hypothesis proposed that thinking is a computational process operating over compositional representations. Complex thoughts comprise simpler components combined through systematic rules, enabling humans to generate infinitely many novel thoughts from finite conceptual primitives. We are not "hapless tourists bound by a limited phrasebook of ideas" (frankland2020concepts), but fluent speakers of an internal, compositional language. However, this view of the mind as a formal logic engine could not explain empirical findings that emerged from psychology laboratories. wason1968reasoning demonstrated a "content effect", where people systematically fail to apply logical rules in an abstract setting but succeed when the identical task is presented as a familiar social rule. tversky1974judgment showed that humans systematically use heuristics and mental shortcuts that violate basic principles of logic and probability theory (ragni2017wason). These results have led to many alternative theoretical accounts. Mental Models Theory proposed that reasoning constructs and manipulates semantic simulations of the world rather than applying syntactic inference rules (johnson1983mental; johnson-laird2010). Dual-Process Theory distinguished fast, intuitive processing from slow, deliberative reasoning (evans2003two; evans2013dual). Bayesian approaches reframed rationality as probabilistic inference under uncertainty, arguing that reasoning updates probabilistic beliefs based on new, ambiguous, and noisy data in ways that approximate Bayesian inference (griffiths2024bayesian; jacobs2011bayesian; oaksford2009precis).
@@
This theoretical diversity in cognitive science, while valuable, has contributed to fragmentation in LLM reasoning research. gandhi2025cognitive characterizes reasoning through verification, subgoal setting, backtracking, and backward chaining, while lee2024reasoning focuses on logical coherence, compositionality, and productivity. Models appear capable under one framework yet fail under another, and researchers lack a shared vocabulary for diagnosing where and why reasoning breaks down. We address this by organizing cognitive theories using Marr’s (marr1982vision) levels of analysis and proposing a unified framework for cognitive elements. The computational level defines the goal of the system: what is being computed and why. The algorithmic and representational level specifies the process: what representations are used for the input/ouput and what algorithms transform these representations to achieve the computational goals. The implementation level concerns the physical realization about how representations and algorithms are instantiated in neural hardware or silicon. For our purposes, we focus on the first two levels. At the computational level, reasoning must satisfy certain fundamental properties to maintain consistency, resolve contradictions, and combine elements compositionally. At the algorithmic level, these goals are realized through specific knowledge structures and the processes that manipulate them (peebles2015thirty; krafft2018levels).
@@
Our proposed framework for cognitive foundations consists of four dimensions (detailed in Table 2). 1. Reasoning invariants specify computational goals—properties that must hold for valid reasoning. 2. Meta-cognitive controls select and monitor processes, determining which strategies to deploy and when to adapt. 3. Reasoning representations describe how knowledge is organized. 4. Reasoning operations specify procedures that construct, evaluate, and transform representations. These dimensions span 28 cognitive elements, providing vocabulary for diagnosing which aspects of human reasoning manifest in language models.
@@
A familiar scenario illustrates how these dimensions interact: a child playing with a pile of LEGO® bricks and building something novel, like a "spaceship" (Figure 1). This creative process is an example of the reasoning invariants of compositionality and productivity, where the goal is to generate a complex, new structure (the spaceship) by combining a finite set of simple components (a few different types of toy bricks). This computational goal is managed by meta-reasoning controls. The child must first decide on a strategy: "Should I have a rough plan first, or go directly into building?" "Should I start with the main body or the wings?" Throughout the process, they must constantly monitor their progress and evaluate their spaceship. For instance, they might observe that the wings don’t look sturdy enough, leading them to reinforce the structure with more blocks. This entire cycle of planning, building, and evaluation relies on the child forming and manipulating a dynamic 3D mental model of the target spaceship, especially the spatial relationships between components. This mental model shifts from a vague concept to a concrete plan through a series of operations. For instance, the child would try out different combinations of pieces, constantly assembling, testing, and then backtracking to refine their final design.
@@
This LEGO example illustrates that reasoning arises from the dynamic interplay of all four dimensions, because the child’s compositional goal requires strategic choices about how to build, which determine what representations to construct, which are then transformed through coordinated operations. Current research tends to isolate individual behaviors—studying verification independently from representation selection, or decomposition separately from meta-cognitive monitoring. However, effective reasoning requires these dimensions to work in concert. The computational drives and constraints from invariants are managed by meta-cognitive controls, which in turn direct the algorithmic-level shifts in representation and the operations that enact them. No single dimension suffices. In the following subsections, we will unpack each dimension in detail, drawing on research in cognitive science to explain what each dimension contributes and examining how each one manifests in current language models.
2.1 Reasoning Invariants: Computational Constraints

### v2

@@ Abstract
Large language models (LLMs) solve complex problems yet fail on simpler variants, suggesting they achieve correct outputs through mechanisms fundamentally different from human reasoning. To understand this gap, we synthesize cognitive science research into a taxonomy of 28 cognitive elements spanning reasoning invariants, meta-cognitive controls, representations for organizing reasoning & knowledge, and transformation operations. We introduce a fine-grained evaluation framework and conduct the first large-scale empirical analysis of 192K traces from 18 models across text, vision, and audio, complemented by 54 human think-aloud traces, which we make publicly available. We find that models under-utilize cognitive elements correlated with success, narrowing to rigid sequential processing on ill-structured problems where diverse representations and meta-cognitive monitoring are critical. Human traces show more abstraction and conceptual processing, while models default to surface-level enumeration. Meta-analysis of 1.6K LLM reasoning papers reveals the research community concentrates on easily quantifiable elements (sequential organization: 55%, decomposition: 60%) but neglecting meta-cognitive controls (self-awareness: 16%) that correlate with success. Models possess behavioral repertoires associated with success but fail to deploy them spontaneously. Leveraging these patterns, we develop test-time reasoning guidance that automatically scaffold successful structures, improving performance by up to 66.7% on complex problems. By establishing a shared vocabulary between cognitive science and LLM research, our framework enables systematic diagnosis of reasoning failures and principled development of models that reason through robust cognitive mechanisms rather than spurious shortcuts, while providing tools to test theories of human cognition at scale.
@@ 1 Introduction
Humans are capable of extrapolating from their existing knowledge to unfamiliar scenarios and generating new knowledge—a process that constitutes reasoning (lombrozo2024learning). In contrast, large language models (LLMs) show failures of generalization that are unintuitive from the standpoint of human reasoning: they master complex skills (chervonyi2025gold; jiang2024survey) while lacking simpler prerequisite ones (mccoy2024embers; mancoridis2025potemkin; dziri2023faith; berglund2023reversal), and solve challenging problems while failing on trivial variants (shao2025spurious; li2025personalized). This dissociation between high benchmark performance and the lack of generalization suggests that models may be achieving correct outputs through mechanisms fundamentally different from the robust cognitive structures underlying human reasoning. Current training and evaluation paradigms reward reasoning outcomes without examining the cognitive processes that produce them (lambert2024tulu3), and therefore cannot distinguish between genuine reasoning and memorization (wu2025reasoning). This creates a measurement crisis: we lack the conceptual framework to characterize what cognitive elements should manifest in models and the empirical methods to assess whether they do.
@@
To understand how cognitive elements manifest in reasoning behaviors, consider a child playing with LEGO blocks. Figure  1 illustrates how the child orchestrates multiple cognitive elements: establishing goals, decomposing the task into parts, verifying connections, monitoring progress, and reformulating failed approaches. This flexible coordination of diverse cognitive processes characterizes reasoning. However, contemporary research on LLM reasoning only attempts scattered investigations of specific behaviors such as verification (gandhi2025cognitive), decomposition (xu2025large), and self-monitoring (marjanovic2025deepseek). Our meta-analysis of 1,598 arXiv LLM reasoning papers on LLM reasoning reveals this fragmentation quantitatively (Figure 2): 55% examine sequential organization (e.g., reasoning step-by-step), while only 8% study backward chaining (e.g., reverse engineering the solution from the answer). This concentration on easily quantifiable behaviors creates a striking asymmetry when compared with human reasoning research, which encompasses diverse phenomena, including analogical transfer (gentner1983structure; holyoak1989analogical), causal inference (sloman2009causal; gopnik2004theory), representational flexibility (ohlsson1992information; knoblich1999constraint), and meta-cognitive control (nelson1990metamemory; flavell1979metacognition). Each of these captures different facets of how humans navigate uncertainty, integrate knowledge, and adapt to novel situations. Without a comprehensive framework spanning this range of cognitive elements, we risk optimizing what we measure rather than what matters, potentially mistaking narrow competence for broad reasoning capability.
@@
Figure 1: An example of cognitive elements present in a reasoning trace for building a LEGO spaceship. We characterize the different elements along the four dimensions of our taxonomy, as shown in Table 2.
@@
Figure 2: Distribution of cognitive element presence across 1,598 arXiv LLM Reasoning papers. Partially present indicates that there is evidence that the element was considered in the design (motivation, method, evaluation) of the paper but was not the primary focus. Present indicates that there is evidence that the element was a conscious and significant design decision. Details provided in Section 5.
@@
To address this gap, we introduce a unified taxonomy of cognitive foundations by synthesizing established theories of problem-solving, mental representation, and meta-cognition (sweller1988cognitive; johnson1983mental; fodor1975language). The taxonomy includes four dimensions: 1) Reasoning invariants capture fundamental properties that must hold for valid reasoning, including logical coherence (not simultaneously believing “the design is stable” and “the design will collapse”), compositionality (understanding “red LEGO cockpit” by combining concepts of color, material, and function). 2) Meta-cognitive controls encompass the executive functions that select and monitor reasoning strategies, such as recognizing when you lack necessary pieces (self-awareness), deciding whether to plan the entire design upfront versus building exploratively (strategy selection). 3) Reasoning representations describe how reasoning and knowledge are organized: hierarchically (structurally decomposing “spaceship” into “body,” “wings,” “cockpit”), causally (conceptually understanding that inadequate support causes structural collapse), or spatially (conceptually tracking how pieces connect in 3D space). 4) Reasoning operations specify the procedures that construct and transform these representations, such as verifying each connection works, backtracking when a wing design fails, or abstracting the principle that heavier components need more support points. Organizing these theories through Marr’s (marr1982vision) levels of analysis (focusing on computational and algorithmic levels) yields 28 specific cognitive elements spanning the space of human reasoning capabilities (Table 2). This taxonomy represents a theoretical contribution: we synthesize existing cognitive science literature into an analytical vocabulary for studying machine reasoning, providing a bridge between human cognitive research and LLM evaluation.
@@
With this taxonomy, we conduct the first large-scale empirical comparison of cognitive elements in human versus LLM reasoning across diverse problem types (jonassen2000toward). We analyze 170K reasoning traces from 18 models in text, vision, and audio modalities, complemented by 54 human think-aloud reasoning traces (Table 2). Using fine-grained span-level annotation validated by human evaluation, we identify which of the 28 cognitive elements appear in each trace, where they occur, and how they are sequenced in thought. Our analysis reveals fundamental differences in reasoning presence and structures. We apply our analysis framework to distinguish between elements that are frequently exhibited by models and those that are most strongly correlated with correct outcomes. Through our analysis, we reveal that models consistently employ cognitive elements that are not the most conducive to success. Specifically, as problems become more ill-defined and non-verifiable (e.g., open-ended dilemmas, diagnosing and solving multifaceted problems), models narrow their selection of elements to rigid strategies despite a broader, more diverse usage of cognitive elements being empirically shown as more effective. Furthermore, we devise a novel reasoning structure representation to encode the structure of cognitive elements within reasoning traces. We find that models frequently choose reasoning structures different from successful structures. Finally, we apply our analysis framework to compare humans and LLMs. We observe that humans employ more frequent abstract and conceptual-level processing, while LLMs rely more heavily on shallow sequential forward chaining with limited corrective structures based on the type and structure of the problem.
@@
Leveraging these behavioral correlations, we introduce test-time reasoning guidance as a targeted intervention to explicitly scaffold cognitive patterns predictive of reasoning success. For instance, on diagnosis problems where successful traces exhibit the behavioral sequence of strategy selection, followed by conceptual processing and causal organization (e.g., “how should I approach diagnosing the issue?” $\rightarrow$ “what are the conceptual factors which could be causing the issue?”), we prompt the model to follow this structure. Test-time reasoning guidance improves performance by up to 66.7% on ill-structured problems while maintaining baseline performance on well-structured ones. This improvement confirms that models possess reasoning capabilities not spontaneously expressed, and that understanding cognitive behavioral patterns can inform more effective model interaction strategies.
@@
This work establishes a taxonomy of cognitive foundations that bridges human reasoning research and LLM evaluation, providing a unified vocabulary for characterizing reasoning processes beyond performance metrics. By synthesizing cognitive science theories through Marr’s levels of analysis, we identify 28 cognitive elements spanning reasoning properties & goals, meta-cognitive controls, reasoning representations, and transformation operations. We conduct the first large-scale empirical analysis in 192K model and human reasoning traces [^1] Publicly released at https://huggingface.co/collections/stellalisy/cognitive-foundations, revealing structural failures and differences from human reasoning. The proposed test-time reasoning guidance that automatically scaffolds successful reasoning structures improves performance by up to 66.7% on complex problems, confirming that models possess latent capabilities that lead to success but fail to deploy them adaptively without explicit structural guidance. Our analysis of 1,598 LLM reasoning papers reveals a critical gap between research emphasis and reasoning requirements. This framework enables the systematic diagnosis of reasoning failures, characterization of how training procedures shape cognitive profiles, and principled development of models that reason through robust cognitive mechanisms rather than spurious shortcuts. Our taxonomy and annotation methodology provide the shared vocabulary necessary for this analysis, grounded in decades of cognitive science research and validated through large-scale empirical study of both human and machine reasoning.
@@ 2 Formalizing Cognitive Foundations for Reasoning
Table 1: Taxonomy of cognitive elements, organized along four main dimensions: Reasoning Invariants, Meta-cognitive Controls, Reasoning Representations, and Reasoning Operations.
@@
Modification heightAdaptive detail management Adjust granularity based on task demands (Rosch1978-ROSPOC-8).
@@
Understanding the extent to which LLMs reason requires a theory of what reasoning is. Without principled criteria distinguishing reasoning from pattern matching, evaluation isn’t possible. Decades of research in cognitive science on problem-solving, mental representation, and meta-cognition provide a powerful starting point. The cognitive revolution of the 1950s introduced a framework of the mind as an information-processing system. Fodor’s (fodor1975language) Language of Thought Hypothesis proposed that thinking is a computational process operating over compositional representations. Complex thoughts comprise simpler components combined through systematic rules, enabling humans to generate infinitely many novel thoughts from finite conceptual primitives. We are not “hapless tourists bound by a limited phrasebook of ideas” (frankland2020concepts), but fluent speakers of an internal, compositional language. However, this view of the mind as a formal logic engine could not explain empirical findings that emerged from psychology laboratories. wason1968reasoning demonstrated a “content effect,” where people systematically fail to apply logical rules in an abstract setting but succeed when the identical task is presented as a familiar social rule. tversky1974judgment showed that humans systematically use heuristics and mental shortcuts that violate basic principles of logic and probability theory (ragni2017wason). These results have led to many alternative theoretical accounts. Mental Models Theory proposed that reasoning constructs and manipulates semantic simulations of the world rather than applying syntactic inference rules (johnson1983mental; johnson-laird2010). Dual-Process Theory distinguished fast, intuitive processing from slow, deliberative reasoning (evans2003two; evans2013dual). Bayesian approaches reframed rationality as probabilistic inference under uncertainty, arguing that reasoning updates probabilistic beliefs based on new, ambiguous, and noisy data in ways that approximate Bayesian inference (griffiths2024bayesian; jacobs2011bayesian; oaksford2009precis).
@@
This theoretical diversity in cognitive science, while valuable, has contributed to fragmentation in LLM reasoning research. gandhi2025cognitive characterizes reasoning through verification, subgoal setting, backtracking, and backward chaining, while lee2024reasoning focuses on logical coherence, compositionality, and productivity. Models appear capable under one framework yet fail under another, and researchers lack a shared vocabulary for diagnosing where and why reasoning breaks down. We address this by organizing cognitive theories using Marr’s (marr1982vision) levels of analysis and proposing a unified framework for cognitive elements. The computational level defines the goal of the system: what is being computed and why. The algorithmic and representational level specifies the process: what representations are used for the input/output and what algorithms transform these representations to achieve the computational goals. The implementation level concerns the physical realization about how representations and algorithms are instantiated in neural hardware or silicon. For our purposes, we focus on the first two levels. At the computational level, reasoning must satisfy certain fundamental properties to maintain consistency, resolve contradictions, and combine elements compositionally. At the algorithmic level, these goals are realized through specific knowledge structures and the processes that manipulate them (peebles2015thirty; krafft2018levels).
@@
Our proposed framework for cognitive foundations consists of four dimensions (detailed in Table 2). 1. Reasoning invariants specify computational goals-properties that must hold for valid reasoning. 2. Meta-cognitive controls select and monitor processes, determining which strategies to deploy and when to adapt. 3. Reasoning representations describe how knowledge is organized. 4. Reasoning operations specify procedures that construct, evaluate, and transform representations. These dimensions span 28 cognitive elements, providing vocabulary for diagnosing which aspects of human reasoning manifest in LLMs.
@@
A familiar scenario illustrates how these dimensions interact: a child playing with a pile of LEGO bricks and building something novel, like a “spaceship” (Figure 1). This creative process engages the reasoning invariants of compositionality and productivity, where the goal is to generate a complex, new structure (the spaceship) by combining a finite set of simple components (a few types of toy bricks). This computational goal is managed by meta-reasoning controls. The child must first decide on a strategy: “Should I have a rough plan first, or go directly into building?” “Should I start with the main body or the wings?” Throughout the process, they must constantly monitor their progress and evaluate their spaceship. They might observe that the wings don’t look sturdy enough, leading them to reinforce the structure with more blocks. This entire cycle of planning, building, and evaluation relies on the child forming and manipulating a dynamic 3D mental model of the target spaceship. This mental model shifts from a vague concept to a concrete plan through a series of operations. For instance, the child would try out different combinations of pieces, constantly assembling, testing, and then backtracking to refine their final design.
@@
This LEGO example illustrates that reasoning arises from the dynamic interplay of all four dimensions, because the child’s computational goal requires strategic choices about how to build, which determine what representations to construct, which are then transformed through coordinated operations. Current research tends to isolate individual elements—studying verification independently of representation selection, or decomposition separately from meta-cognitive monitoring. However, effective reasoning requires these dimensions to work in concert. The computational drives and constraints from invariants are managed by meta-cognitive controls, which in turn direct the algorithmic-level shifts in representation and the operations that enact them. No single dimension suffices. In the following subsections, we will unpack each dimension in detail, drawing on research in cognitive science to explain what each dimension contributes and examining how they manifests in current LLMs.
2.1 Reasoning Invariants: Properties & Goals